# Short-duration GRB 250221A Afterglow Driven by Two-Component Jets from merger of compact star

### v1

@@ III.4 Physical offset and star-forming rate of the host galaxy
The host galaxies of LGRBs are typically actively star-forming and metal-poor, with burst locations that are closely offset from their host centers (Fruchter et al., 2006). This is consistent with the model in which LGRBs originate from the collapse of massive stars (MacFadyen & Woosley, 1999). On the contrary, SGRBs are generally found in older stellar populations, such as elliptical galaxies with low star formation rates, while some occur in star forming galaxies (Berger et al., 2005; Berger, 2014). Their explosion sites are often offset from the centers of their host galaxies, with physical offsets ranging from a few to several tens of kiloparsecs (Fong & Berger, 2013). The K–S test shows that the physical offset distributions of long and short GRBs do not originate from the same parent population (Li et al., 2016). We compare the physical offset of the host galaxy for GRB 250221A (e.g., $7^{+12}_{-7}$ kpc) with that of other long and short GRB samples observed by the Hubble Space Telescope (see Figure 2(a)) (Fong et al., 2010; Blanchard et al., 2016). It is found that the physical offset of GRB 250221A is larger than that of most LGRBs, but is statistically consistent with that of typical SGRBs. Moreover, we also compare the star formation rate (SFR $\sim 3~M_{\odot}~\rm yr^{-1}$) of the host galaxy for GRB 250221A with those of long (Dong et al., 2025) and short (Li et al., 2016; Nugent et al., 2022) GRBs  [^2] The data of several short GRBs are taken from the GHostS websit: https://www.grbhosts.org/ (see Figure 2(b)). We find that SFR alone cannot unambiguously determine the origin of GRB 250221A.
@@ IV Physical Interpretation
The burst properties and empirical correlations or distributions (duration, spectral hardness, location in the Amati correlation, $\varepsilon$-value, $f_{\rm eff}$ parameter, and physical offset) are in favor of a compact binary merger origin (also see Angulo-Valdez et al. (2026)). However, Angulo-Valdez et al. (2026) adopts energy injection into a jet to interpret the late-time re-brightening in both X-ray and optical bands $\sim 0.6$ days after the burst trigger, and obtains a dense circumburst medium with $n\sim 80\rm~cm^{-3}$. Such a high medium density is inconsistent with the environment of a compact binary merger, but is in favor of the typical environment of a Collapsar (Angulo-Valdez et al., 2026).
In order to reconcile the above contradictions, we invoke a two-component jet model which is first proposed by Jin et al. (2007) within the scenario of a compact binary merger to explain the multiwavelength afterglow of GRB 250221A. The two-component jet consists of an ultra-relativistic narrow jet core surrounded by a mildly relativistic wide jet component, and such a model is widely used to explain the complex evolution of GRB afterglows (Berger et al., 2003; Peng et al., 2005; Granot et al., 2006; Racusin et al., 2008). Within this scenario, the narrow jet is primarily responsible for the prompt $\gamma$-ray and the early afterglow emission, whereas the emission from the wide, slow jet component rises to the late-time re-brightening afterglow.
We employ the $VegasAfterglow$ code which is proposed by Wang et al. (2026) to fit the multiwavelength afterglow data of GRB 250221A. This code describes an axisymmetric jet model, and the two-component jet structure can be expressed as follows:
@@
Here, $E_{\rm K,iso}$, $\rm\Gamma_{0}$ and $\rm\theta_{c}$ represent the isotropic equivalent kinetic energy, initial bulk Lorentz factor, and half-opening angle of the narrow jet core, respectively. $E_{\rm K,iso,w}$, $\rm\Gamma_{0,w}$ and $\rm\theta_{w}$ correspond to the parameters of the wide jet component. By assuming a constant density profile of the ISM for the afterglow, we adopt the Markov Chain Monte Carlo (MCMC) method provided by Vegasafterglow (Wang et al., 2026) to perform the posterior parameter inference [^3] Moreover, other several public codes of afterglow are also proposed in the literatures to fit the afterglow data, such as afterglowpy (Ryan et al., 2020, 2024), jetsimpy (Wang et al., 2024a, 2025), ASGARD (Ren et al., 2024), and redback (Sarin et al., 2024).. The basic parameters, $E_{\rm K,iso}$, $\rm\Gamma_{0}$, $\rm\theta_{c}$, $E_{\rm K,iso,w}$, $\rm\Gamma_{0,w}$, $\rm\theta_{w}$, $n$, $\rm p$, $\epsilon_{\rm e}$, and $\epsilon_{\rm B}$, are shown in Table 1. Of these, the parameters $\rm\theta_{c}$, $\rm\theta_{w}$, and $\rm p$ are sampled on a linear scale, while the remaining parameters are sampled on a logarithmic scale. The fraction of electrons $\xi_{\rm e}=1$ is fixed as commonly adopted in afterglow modeling. The best-fit results for the multiwavelength afterglow light curves are shown in Figure 3, and the model parameters are summarized in Table 1. The corner plots of the posterior density distributions from the MCMC samples are shown in Figure 4.
The fitting results show that the afterglow of GRB 250221A can be explained by a narrow jet core with a half-opening angle of $\rm\theta_{c}\sim 3.8^{\circ}$ and an initial bulk Lorentz factor $\rm\Gamma_{0}\sim 507$, surrounded by a mildly relativistic wide jet with $\rm\theta_{w}\sim 4.4^{\circ}$ and $\rm\Gamma_{0,w}\sim 46$. Within this scenario, the emission from the narrow jet has entered its decay phase at the beginning of the observation of the afterglow, while the emission from the wide jet continues to rise and eventually becomes dominant. It can result in the distinct re-brightening at around 0.6 days observed in the optical and X-ray bands. Within the framework of two-component jet model, the derived circumburst medium density is $n\sim 0.72~\rm cm^{-3}$ which is much lower than that of previous work by adopting an energy-injection scenario (e.g., Angulo-Valdez et al. (2026)) inferred a density around $n\sim 80~\rm cm^{-3}$. Moreover, we compare the obtained medium density with that of other short GRBs taken from Fong et al. (2015) (see right of Figure 3), and find that although this medium density is slightly higher than that of SGRBs in merger environments, it still falls into a reasonable and acceptable range (Fong et al., 2015). Together with the burst properties and empirical correlations or distributions of GRB 250221A, it suggests that the short-duration GRB 250221A originated from the merger of compact stars.
@@ V Possible kilonova emission
Follow-up optical observations of GRB 250221A from the COLIBRI and VLT reveal a light curve showing an initially decay phase and a re-brightening at approximately 0.6 days (Angulo-Valdez et al., 2026). In order to investigate whether this feature could be caused by the possible associated kilonova emission, we calculate the possible kilonova emission in the $i$ and $r$ bands by adopting $M_{\rm ej}=(10^{-3}-10^{-1})\rm~M_{\odot}$, velocity $\beta=0.1$ and opacity $\kappa=1~\rm cm^{2}~g^{-1}$ (Hotokezaka et al., 2013). In our calculation, we only consider the r-process itself, and the equations to describe the evolution of the ejecta dynamics and the internal energy are taken from (Yu et al., 2013) and (Yuan et al., 2021). The results of the numerical calculations and the comparison with observations are shown in Figure 5. We find that the potential kilonova emission associated with GRB 250221A at a redshift of $z=0.768$ would be extremely faint. Even at its peak, the kilonova flux remains two orders of magnitude lower than the observed optical data. Therefore, the re-brightening feature observed in the optical afterglow of GRB 250221A is unlikely to be caused by kilonova emission.
On the other hand, one possible is that the central engine is a magnetar which may be formed after the mergers of binary neutron stars, the spin energy of magnetar can be injected into kilonova to increase its peak luminosity (Yu et al., 2013; Metzger & Piro, 2014; Gao et al., 2017; Wang et al., 2024b; Ai et al., 2025). If this is the case, it requires an extremely spin-down luminosity and timescale of magnetar to power such optical rebrightening for GRB 250221A. However, lack of enough observational evidence to support the magnetar is resided as the central engine of GRB 250221A.
@@ VI Conclusion and Discussion
GRB 250221A is a short-duration burst simultaneously detected by Swift and Konus-Wind with a duration of $T_{90}\sim 1.8$ s at a redshift of $z=0.768$. The prompt $\gamma$-ray emission of this burst consists of a single short pulse, and the CPL model can well fit the spectrum of GRB 250221A with a peak energy $E_{\rm p}=242^{+26}_{-22}$ keV. By adopting the measured redshift $z=0.768$ (Angulo-Valdez et al., 2026), one can estimate the isotropic $\gamma$-ray energy $E_{\rm\gamma,iso}=(2.8\pm 0.2)\times 10^{51}\rm~erg$ (Frederiks et al., 2025). The physical offset and star formation rate of the host galaxy of GRB 250221A are about $7^{+12}_{-7}$ kpc and $\sim 3~M_{\odot}~\rm yr^{-1}$, respectively. More interestingly, both the optical and X-ray afterglow of GRB 250221A exhibit an initial decay followed by a remarkable re-brightening around 0.6 days after the trigger.
@@
By comparing the properties and empirical correlations or distributions (e.g., duration, spectral hardness, location in the Amati correlation, $\varepsilon$-value, $f_{\rm eff}$ parameter, and physical offset) of GRB 250221A with those of other Type I and Type II GRBs, it is found that GRB 250221A shares similar properties and empirical correlations or distributions with Type I GRBs, suggesting that GRB 250221A favors a compact binary merger origin. However, Angulo-Valdez et al. (2026) invoked a model of energy injection into a jet to interpret the late-time re-brightening, and obtained a medium density as high as $n\sim 80\rm~cm^{-3}$ , which is inconsistent with typical merger environments (Fong et al., 2015), but is in favor of a collapsar environment.
@@
Alternatively, in order to reconcile the above contradictions, we propose a two-component jet model within the scenario of compact binary merger to explain the multiwavelength afterglow of GRB 250221A. The narrow jet is primarily responsible for the prompt $\gamma$-ray emission and the early afterglow emission, whereas the emission from the wide, slow jet component rises to the late-time re-brightening afterglow. Within this scenario, one can obtain the relativistic narrow jet ($\rm\theta_{c}\sim 3.8^{\circ}$, $\rm\Gamma_{0}\sim 507$) which produces an initial rapid decay phase after deceleration, and the mildly relativistic wide jet ($\rm\theta_{w}\sim 4.4^{\circ}$, $\rm\Gamma_{0,w}\sim 46$) dominates the late-time emission to result in a pronounced rise in the light curve around 0.6 days after the trigger. On the other hand, one can obtain the medium density $n\sim 0.72~\rm cm^{-3}$, which is much lower than the value calculated in Angulo-Valdez et al. (2026) by invoking the energy injection into a jet. Although such medium density is slightly higher than that of SGRBs in merger environments, it still falls within the reasonable and acceptable range (Fong et al., 2015). In any case, together with the burst properties and empirical correlations or distributions of GRB 250221A, as well as the lower medium density, it suggests that the short-duration GRB 250221A still originates from the mergers of compact stars.
@@
Since the fitting results by adopting the two-component jet model, are dependent on the selected structure of jet. For example, Ren et al. (2024) performed a top-hat inner core encompassed by a power-law-structured outer wing as the structure of the two-component jet. Moreover, the algorithm of different afterglow codes (e.g., afterglowpy, jetsimpy, ASGARD, and redback) can also result in a little difference even for the same data. On the other hand, one needs to clarify that the two-component jet model we adopt is not the only physical model to interpret the observational afterglow data of GRB 250221A, and we cannot rule out other possible physical mechanisms (Yang et al., 2024). However, together with the burst properties and empirical correlations or distributions of GRB 250221A, it provides at least a self-consistent physical explanation of the data.
@@
Within the scenario of the two-component jet model, the problem is that the wide component possesses higher energy than the narrow core, while their half-opening angles differ by only about 0.015 rad. Based on the 3-D simulations from Nathanail et al. (2021), they adopted a continuous angular structure in which both the energy and the Lorentz factor peak at finite angles away from the axis, often surrounded by slower, baryon-rich outflows, and it is invoked to interpret off-axis GRB and its kilonova emission. The two‑component model is fundamentally a parameterized approximation designed to capture strong angular stratification within the jet, rather than to represent two physically distinct jet layers. However, we adopt the two-component picture of a narrow-fast core encircled by a wide-slower component, and it is quite different from the picture in Nathanail et al. (2021).
@@
Moreover, we also calculate the theoretical kilonova emission that is possibly associated with GRB 250221A in the $i$- and $r$-bands. It is found that the potential kilonova emission associated with GRB 250221A at a redshift of $z=0.768$ would be extremely faint. Even at its peak, the kilonova emission remains two orders of magnitude lower than the observed optical data. Therefore, the re-brightening feature observed in the optical afterglow of GRB 250221A is unlikely to be caused by the kilonova emission.

### v2

@@ III.4 Physical offset and star-forming rate of the host galaxy
The host galaxies of LGRBs are typically actively star-forming and metal-poor, with burst locations that are closely offset from their host centers (Fruchter et al., 2006). This is consistent with the model in which LGRBs originate from the collapse of massive stars (MacFadyen & Woosley, 1999). On the contrary, SGRBs are generally found in older stellar populations, such as elliptical galaxies with low SFRs, while some occur in star forming galaxies (Berger et al., 2005; Berger, 2014). Their explosion sites are often offset from the centers of their host galaxies, with physical offsets ranging from a few to several tens of kiloparsecs (Fong & Berger, 2013). The Kolmogorov–Smirnov test shows that the physical offset distributions of long and short GRBs do not originate from the same parent population (Li et al., 2016). We compare the physical offset of the host galaxy for GRB 250221A (e.g., $7^{+12}_{-7}$ kpc) with that of other long and short GRB samples observed by the Hubble Space Telescope (see Figure 2(a)) (Fong et al., 2010; Blanchard et al., 2016). It is found that the physical offset of GRB 250221A is larger than that of most LGRBs, but is statistically consistent with that of typical SGRBs. Moreover, we also compare the SFR (SFR $\sim 3~M_{\odot}~\rm yr^{-1}$) of the host galaxy for GRB 250221A with those of long (Dong et al., 2025) and short (Li et al., 2016; Nugent et al., 2022) GRBs  [^2] The data of several short GRBs are taken from the GHostS websit: https://www.grbhosts.org/ (see Figure 2(b)). We find that SFR alone cannot unambiguously determine the origin of GRB 250221A.
@@ IV Physical Interpretation
The burst properties and empirical correlations or distributions (duration, spectral hardness, location in the Amati correlation, $\varepsilon$ value, $f_{\rm eff}$ parameter, and physical offset) are in favor of a compact binary merger origin (also see Angulo-Valdez et al. (2026)). However, Angulo-Valdez et al. (2026) adopt energy injection into a jet to interpret the late-time rebrightening in both X-ray and optical bands $\sim 0.6$ days after the burst trigger, and obtain a dense circumburst medium with $n\sim 80\rm~cm^{-3}$. Such a high medium density is inconsistent with the environment of a compact binary merger, but is in favor of the typical environment of a collapsar (Angulo-Valdez et al., 2026).
In order to reconcile the above contradictions, we invoke a two-component jet model, which was first proposed by Jin et al. (2007) within the scenario of a compact binary merger to explain the multiwavelength afterglow of GRB 250221A. The two-component jet consists of an ultrarelativistic narrow jet core surrounded by a mildly relativistic wide jet component, and such a model is widely used to explain the complex evolution of GRB afterglows (Berger et al., 2003; Peng et al., 2005; Granot et al., 2006; Racusin et al., 2008). Within this scenario, the narrow jet is primarily responsible for the prompt gamma-ray and the early afterglow emission, whereas the emission from the wide, slow jet component rises to the late-time rebrightening afterglow.
We employ the $VegasAfterglow$ code, which is proposed by Wang et al. (2026) to fit the multiwavelength afterglow data of GRB 250221A. This code describes an axisymmetric jet model, and the two-component jet structure can be expressed as follows:
@@
Here, $E_{\rm K,iso}$, $\rm\Gamma_{0}$ and $\rm\theta_{c}$ represent the isotropic equivalent kinetic energy, initial bulk Lorentz factor, and half-opening angle of the narrow jet core, respectively. $E_{\rm K,iso,w}$, $\rm\Gamma_{0,w}$ and $\rm\theta_{w}$ correspond to the parameters of the wide jet component. By assuming a constant density profile of the ISM for the afterglow, we adopt the Markov Chain Monte Carlo (MCMC) method provided by Vegasafterglow (Wang et al., 2026) to perform the posterior parameter inference [^3] Moreover, several other public codes of afterglow are also proposed in the literature to fit the afterglow data, such as afterglowpy (Ryan et al., 2020, 2024), jetsimpy (Wang et al., 2024a, 2025), ASGARD (Ren et al., 2024), and redback (Sarin et al., 2024).. The basic parameters, $E_{\rm K,iso}$, $\rm\Gamma_{0}$, $\rm\theta_{c}$, $E_{\rm K,iso,w}$, $\rm\Gamma_{0,w}$, $\rm\theta_{w}$, $n$, $p$, $\epsilon_{\rm e}$, and $\epsilon_{\rm B}$, are shown in Table 1. Of these, the parameters $\rm\theta_{c}$, $\rm\theta_{w}$, and $p$ are sampled on a linear scale, while the remaining parameters are sampled on a logarithmic scale. The fraction of electrons $\xi_{\rm e}=1$ is fixed as commonly adopted in afterglow modeling. The best-fit results for the multiwavelength afterglow light curves are shown in Figure 3, and the model parameters are summarized in Table 1. The corner plots of the posterior density distributions from the MCMC samples are shown in Figure 4.
The fitting results show that the afterglow of GRB 250221A can be explained by a narrow jet core with a half-opening angle of $\rm\theta_{c}\sim 3.8^{\circ}$ and an initial bulk Lorentz factor $\rm\Gamma_{0}\sim 507$, surrounded by a mildly relativistic wide jet with $\rm\theta_{w}\sim 4.4^{\circ}$ and $\rm\Gamma_{0,w}\sim 46$. Within this scenario, the emission from the narrow jet has entered its decay phase at the beginning of the observation of the afterglow, while the emission from the wide jet continues to rise and eventually becomes dominant. It can result in the distinct rebrightening at around 0.6 days observed in the optical and X-ray bands. Within the framework of the two-component jet model, the derived circumburst medium density is $n\sim 0.72~\rm cm^{-3}$, which is much lower than that of previous work by adopting an energy-injection scenario (e.g., Angulo-Valdez et al. (2026)) inferred a density around $n\sim 80~\rm cm^{-3}$. Moreover, we compare the obtained medium density with that of other short GRBs taken from Fong et al. (2015) (see right of Figure 3), and find that although this medium density is slightly higher than that of SGRBs in merger environments, it still falls into a reasonable and acceptable range (Fong et al., 2015). Together with the burst properties and empirical correlations or distributions of GRB 250221A, it suggests that the SGRB 250221A originated from the merger of compact stars.
@@ V Possible kilonova emission
Follow-up optical observations of GRB 250221A from the COLIBRI and VLT reveal a light curve showing an initial decay phase and a rebrightening at approximately 0.6 days (Angulo-Valdez et al., 2026). In order to investigate whether this feature could be caused by the possible associated kilonova emission, we calculate the possible kilonova emission in the $i$ and $r$ bands by adopting $M_{\rm ej}=(10^{-3}-10^{-1})\rm~M_{\odot}$, velocity $\beta=0.1$ and opacity $\kappa=1~\rm cm^{2}~g^{-1}$ (Hotokezaka et al., 2013). In our calculation, we only consider the r-process itself, and the equations to describe the evolution of the ejecta dynamics and the internal energy are taken from (Yu et al., 2013) and (Yuan et al., 2021). The results of the numerical calculations and the comparison with observations are shown in Figure 5. We find that the potential kilonova emission associated with GRB 250221A at a redshift of $z=0.768$ would be extremely faint. Even at its peak, the kilonova flux remains 2 orders of magnitude lower than the observed optical data. Therefore, the rebrightening feature observed in the optical afterglow of GRB 250221A is unlikely to be caused by kilonova emission.
On the other hand, one possibility is that the central engine is a magnetar, which may be formed after the mergers of binary neutron stars, and the spin energy of magnetar can be injected into kilonova to increase its peak luminosity (Yu et al., 2013; Metzger & Piro, 2014; Gao et al., 2017; Wang et al., 2024b; Ai et al., 2025). If this is the case, it requires an extreme spin-down luminosity and timescale of a magnetar to power such optical rebrightening for GRB 250221A. However, there is a lack of enough observational evidence to support that the magnetar resides as the central engine of GRB 250221A.
@@ VI Conclusion and Discussion
GRB 250221A is a short-duration burst simultaneously detected by Swift and Konus-Wind with a duration of $T_{90}\sim 1.8$ s at a redshift of $z=0.768$. The prompt gamma-ray emission of this burst consists of a single short pulse, and the CPL model can well fit the spectrum of GRB 250221A with a peak energy $E_{\rm p}=242^{+26}_{-22}$ keV. By adopting the measured redshift $z=0.768$ (Angulo-Valdez et al., 2026), one can estimate the isotropic gamma-ray energy $E_{\rm\gamma,iso}=(2.8\pm 0.2)\times 10^{51}\rm~erg$ (Frederiks et al., 2025). The physical offset and SFR of the host galaxy of GRB 250221A are about $7^{+12}_{-7}$ kpc and $\sim 3~M_{\odot}~\rm yr^{-1}$, respectively. More interestingly, both the optical and X-ray afterglow of GRB 250221A exhibit an initial decay followed by a remarkable rebrightening around 0.6 days after the trigger.
@@
By comparing the properties and empirical correlations or distributions (e.g., duration, spectral hardness, location in the Amati correlation, $\varepsilon$ value, $f_{\rm eff}$ parameter, and physical offset) of GRB 250221A with those of other Type I and Type II GRBs, it is found that GRB 250221A shares similar properties and empirical correlations or distributions with Type I GRBs, suggesting that GRB 250221A favors a compact binary merger origin. However, Angulo-Valdez et al. (2026) invoked a model of energy injection into a jet to interpret the late-time rebrightening, and obtained a medium density as high as $n\sim 80\rm~cm^{-3}$ , which is inconsistent with typical merger environments (Fong et al., 2015), but is in favor of a collapsar environment.
@@
Alternatively, in order to reconcile the above contradictions, we propose a two-component jet model within the scenario of compact binary merger to explain the multiwavelength afterglow of GRB 250221A. The narrow jet is primarily responsible for the prompt gamma-ray emission and the early afterglow emission, whereas the emission from the wide, slow jet component rises to the late-time rebrightening afterglow. Within this scenario, one can obtain the relativistic narrow jet ($\rm\theta_{c}\sim 3.8^{\circ}$, $\rm\Gamma_{0}\sim 507$) which produces an initial rapid decay phase after deceleration, and the mildly relativistic wide jet ($\rm\theta_{w}\sim 4.4^{\circ}$, $\rm\Gamma_{0,w}\sim 46$) dominates the late-time emission to result in a pronounced rise in the light curve around 0.6 days after the trigger. On the other hand, one can obtain the medium density $n\sim 0.72~\rm cm^{-3}$, which is much lower than the value calculated in Angulo-Valdez et al. (2026) by invoking the energy injection into a jet. Although such medium density is slightly higher than that of SGRBs in merger environments, it still falls within the reasonable and acceptable range (Fong et al., 2015). In any case, together with the burst properties and empirical correlations or distributions of GRB 250221A, as well as the lower medium density, it suggests that the SGRB 250221A still originates from the mergers of compact stars.
@@
The fitting results by adopting the two-component jet model are dependent on the selected structure of the jet. For example, Ren et al. (2024) performed a top-hat inner core encompassed by a power-law-structured outer wing as the structure of the two-component jet. Moreover, the algorithm of different afterglow codes (e.g., afterglowpy, jetsimpy, ASGARD, and redback) can also result in a little difference even for the same data. On the other hand, one needs to clarify that the two-component jet model we adopt is not the only physical model to interpret the observational afterglow data of GRB 250221A, and we cannot rule out other possible physical mechanisms (Yang et al., 2024). However, together with the burst properties and empirical correlations or distributions of GRB 250221A, it provides at least a self-consistent physical explanation of the data.
@@
Within the scenario of the two-component jet model, the problem is that the wide component possesses higher energy than the narrow core, while their half-opening angles differ by only about 0.01 rad. Based on the 3-D simulations from Nathanail et al. (2021), they adopted a continuous angular structure in which both the energy and the Lorentz factor peak at finite angles away from the axis, often surrounded by slower, baryon-rich outflows, and it is invoked to interpret off-axis GRB and its kilonova emission. The two‑component model is fundamentally a parameterized approximation designed to capture strong angular stratification within the jet, rather than to represent two physically distinct jet layers. However, we adopt the two-component picture of a narrow-fast core encircled by a wide-slower component, and it is quite different from the picture in Nathanail et al. (2021).
@@
Moreover, we also calculate the theoretical kilonova emission that is possibly associated with GRB 250221A in the $i$- and $r$ bands. It is found that the potential kilonova emission associated with GRB 250221A at a redshift of $z=0.768$ would be extremely faint. Even at its peak, the kilonova emission remains two orders of magnitude lower than the observed optical data. Therefore, the rebrightening feature observed in the optical afterglow of GRB 250221A is unlikely to be caused by the kilonova emission.